OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ x\equiv 3(mod~5)\hspace{.33em}\\~\\ x\equiv 5(mod~7)\hspace{.33em}\\~\\ x\equiv 7(mod~11)\hspace{.33em}\\~\\ }\end{align}\) find \(x\).

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{gcd(5,7)=\gcd(5,11)=\gcd(11,7)=1\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (kainui):

I don't know if this is technically how you're expected to solve it since I know there's like a chinese remainder theorem or something for this but I couldn't remember it, so I am going to try to figure it out like this: \[\Large x=5a+3 \\ \Large x=7b+5 \\ \Large x=11c+7\] From here you can equate all of the x's in 3 different ways like this: \[\Large 5a+3=7b+5 \\ \Large 5a+3=11c+7 \\ \Large 7b+5=11c+7\] Then rearrange them like this: \[\Large 5a-7b=2 \\ \Large 5a-11c=4 \\ \Large 7b-11c=2\] Plug it into a 3x3 matrix like this: \[\Large \left[\begin{matrix}5 & -7 & 0 \\ 5 & 0&-11 \\ 0 & 7 & -11\end{matrix}\right] \left(\begin{matrix}a \\ b \\ c\end{matrix}\right) = \left(\begin{matrix}2 \\ 4 \\ 2\end{matrix}\right)\] Multiply by the inverse... And wait, there is no inverse since the determinant is 0 if you look down the antidiagonal... :,( Nevermind good luck, I'll keep tring.

OpenStudy (mathmath333):

yes the chinese remainder

OpenStudy (mathmath333):

but i there must be an easy way without matrix

ganeshie8 (ganeshie8):

You want the remainders to be 3,5,7 when you divide by 5,7,11 respectively. One way to work it is by trial and error : \(x \equiv 3\pmod 5\) : 3, 8, 13, 18, ... \(x \equiv 5\pmod 7\) : 5, 12, 19, 26, 33, ... \(x \equiv 7\pmod {11}\) : 7, 18, 29, 40,... keep going till you find a common term among all 3 sequences

ganeshie8 (ganeshie8):

chinese remainder thm gives you answer in one line, but lets see another easy way to work this system (similar to matrix method but w/o using matrices directly)

ganeshie8 (ganeshie8):

Look at first congruence : \(x \equiv 3\pmod 5\) that means \(x = 3 + 5k\) where \(k\) is an integer substitute this in second congruence

ganeshie8 (ganeshie8):

\(x \equiv 5 \pmod 7\) \(3+5k \equiv 5 \pmod 7\) solve \(k\)

ganeshie8 (ganeshie8):

we want an integer solution

OpenStudy (mathmath333):

oh wait \(k=-1\)

ganeshie8 (ganeshie8):

\(x \equiv 5 \pmod 7\) \(3+5k \equiv 5 \pmod 7\) subtracting 3 both sides you get \(5k \equiv 2 \pmod 7\) now notice that k = 6 satisfies this congruence so the solution is \(k \equiv 6 \pmod 7\) which is same as \(k = 7m+6\) where \(m\) is any integer

ganeshie8 (ganeshie8):

yes! k = -1 is same as 6 in mod 7

ganeshie8 (ganeshie8):

So far we have this : x = 3 + 5k k = 7m + 6 combining you get x = 33 + 5*7m plug this in third congruence and solve m

ganeshie8 (ganeshie8):

\(x\equiv 7 \pmod {11}\) \(33 + 5\cdot 7 m \equiv 7 \pmod {11}\) solve \(m\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ 33+35m=7+11n \hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

it will be easy to work it using mod properties

OpenStudy (mathmath333):

is the last second paragraph above did u mean \(\large 11\) divides 33 so the remainder would be 0 :

ganeshie8 (ganeshie8):

Yes ! sorry it was a typo let me fix it :)

ganeshie8 (ganeshie8):

\(x\equiv 7 \pmod {11}\) \(33 + 5\cdot 7 m \equiv 7 \pmod {11}\) \(\color{Red}{11}\) divides \(33\) so the remainder would be \(0\) : \(0+ 5\cdot 7 m \equiv 7 \pmod {11}\) \(35 m \equiv 7 \pmod {11}\)

ganeshie8 (ganeshie8):

35 leaves a remainder of 2 when divide by 11, so \(35 m \equiv 7 \pmod {11}\) simplifies to \(2 m \equiv 7 \pmod {11}\) see what value of \(m\) satisfies this congruence

OpenStudy (mathmath333):

m should be integer ?

ganeshie8 (ganeshie8):

in number theory we only deal with integers

ganeshie8 (ganeshie8):

unless otherwise explicitly stated all variables are assumed to be integers :)

ganeshie8 (ganeshie8):

Notice when you're working in mod11, you just have 11 values to test : {0,1,2,..,10}

OpenStudy (mathmath333):

oh \(m=9\) ?

ganeshie8 (ganeshie8):

Right!

ganeshie8 (ganeshie8):

since you're in mod11, the solution is m = 9 + 11n where n is ANY/ALL integers

ganeshie8 (ganeshie8):

plug that in previous equation for x and you're done!

ganeshie8 (ganeshie8):

``` So far we have this : x = 3 + 5k k = 7m + 6 combining you get x = 33 + 5*7m plug this in third congruence and solve m ```

ganeshie8 (ganeshie8):

plugin m = 9 + 11n above ^

ganeshie8 (ganeshie8):

x = 33 + 5*7m x = 33 + 5*7(9+11n) simplifying you get x = 348 + 5*7*11n thats the final unique incongruent solution in mod 5*7*11

OpenStudy (mathmath333):

x=348

ganeshie8 (ganeshie8):

notice 348 is the first positive number that satisfies all the 3 given congruences

OpenStudy (mathmath333):

thank u both

ganeshie8 (ganeshie8):

chinese remainder thm is just a compressed version of our previous method : \[x \equiv 3\pmod 5\\x\equiv 5\pmod 7\\x\equiv 7 \pmod {11}\] As promised chinese remainder thm gives you answer in one line \[x \equiv 3 (7\cdot 11)[ (7\cdot 11 )^{-1}\pmod 5] \\+\\5 (5\cdot 11)[ (5\cdot 11)^{-1}\pmod 7] \\+ \\7 (5\cdot 7)[ (5\cdot 7)^{-1}\pmod {11}]\\ \pmod{5\cdot 7 \cdot 11} \] you shoudl get 348 after working it :P

OpenStudy (mathmath333):

but i dont get how will u find the mod of inverse \(\large \begin{align} \color{black}{ (7\cdot 11)^{-1} (mod~5)\hspace{.33em}\\~\\ }\end{align}\)

ganeshie8 (ganeshie8):

that notation refers to "multiplicative inverse"

ganeshie8 (ganeshie8):

finding \((7\cdot 11)^{-1}\) is same as solving the congruence : \[7\cdot 11 x \equiv 1 \pmod 5\]

ganeshie8 (ganeshie8):

its basically asking you for the value of x that produces 1 when multiplied by 7.11 in mod 5

ganeshie8 (ganeshie8):

consider a simpler example : \(2^{-1} \pmod {5}\)

ganeshie8 (ganeshie8):

To find \(2^{-1} \pmod {5}\), we solve \[ 2x\equiv 1 \pmod {5}\]

OpenStudy (mathmath333):

x=3 ?

ganeshie8 (ganeshie8):

Yes! since x = 3 satisfues above congruence, the multiplciative inverse of 2 in mod 5 is 3 : \[2^{-1} \pmod{5} = 3\]

ganeshie8 (ganeshie8):

maybe lets work the previous chinese monster expression to get more practice

ganeshie8 (ganeshie8):

\[x \equiv 3 (7\cdot 11)[ \color{Red}{(7\cdot 11 )^{-1}\pmod 5}] \\+\\5 (5\cdot 11)[ (5\cdot 11)^{-1}\pmod 7] \\+ \\7 (5\cdot 7)[ (5\cdot 7)^{-1}\pmod {11}]\\ \pmod{5\cdot 7 \cdot 11}\] work that red part first

ganeshie8 (ganeshie8):

finding \((7\cdot 11)^{-1}\) is same as solving the congruence : \[7 \cdot 11x \equiv 1\pmod{5}\] find a \(x\) that satisfies above congruence

OpenStudy (mathmath333):

\(x=231\)

ganeshie8 (ganeshie8):

you're in mod5 so you can play with smaller numbers : 0,1,2,3,4

OpenStudy (mathmath333):

i mean that red part value is 231

ganeshie8 (ganeshie8):

\[7 \cdot 11x \equiv 1\pmod{5}\] notice \(7\equiv 2\) and \(11\equiv 1\) in mod 5, so : \[2 \cdot 1x \equiv 1\pmod{5}\]

OpenStudy (mathmath333):

x=3

ganeshie8 (ganeshie8):

yes so the red part evaluates to 3

OpenStudy (mathmath333):

ohh

ganeshie8 (ganeshie8):

\[\color{Red}{(7\cdot 11 )^{-1}\pmod 5} = 3\]

ganeshie8 (ganeshie8):

similarly work other two inverses also : \[x \equiv 3 (7\cdot 11)[ \color{Red}{(7\cdot 11 )^{-1}\pmod 5}] \\+\\5 (5\cdot 11)[ \color{Red}{(5\cdot 11)^{-1}\pmod 7}] \\+ \\7 (5\cdot 7)[\color{Red}{ (5\cdot 7)^{-1}\pmod {11}}]\\ \pmod{5\cdot 7 \cdot 11}\]

ganeshie8 (ganeshie8):

once you have the inverse you just plug them in and simplify you will end up with 348

OpenStudy (mathmath333):

next 2nd red part =6

ganeshie8 (ganeshie8):

Right!

OpenStudy (mathmath333):

next 4

ganeshie8 (ganeshie8):

i think i got 6, one sec..

OpenStudy (mathmath333):

yes 6 sry

ganeshie8 (ganeshie8):

yes! next plug and chug

ganeshie8 (ganeshie8):

\[x \equiv 3 (7\cdot 11)[ \color{Red}{3}] \\+\\5 (5\cdot 11)[ \color{Red}{6}] \\+ \\7 (5\cdot 7)[\color{Red}{ 6}]\\ \pmod{5\cdot 7 \cdot 11}\]

OpenStudy (mathmath333):

that is like 3813

ganeshie8 (ganeshie8):

Correct! 3813 in mod 5*7*11

ganeshie8 (ganeshie8):

reduce it

OpenStudy (mathmath333):

lol there must a short way

ganeshie8 (ganeshie8):

short way for what

OpenStudy (mathmath333):

for evaluating \(x \equiv 3 (7\cdot 11)[ \color{Red}{3}] \\+\\5 (5\cdot 11)[ \color{Red}{6}] \\+ \\7 (5\cdot 7)[\color{Red}{ 6}]\\ \pmod{5\cdot 7 \cdot 11}\)

ganeshie8 (ganeshie8):

ofcourse you can use mod properties for each sum

ganeshie8 (ganeshie8):

\[3(7*11)(3) = 693 \equiv -77 \pmod {5*7*11}\] but its not so much important how you work it.. your goal is just to get a positive integer less than 5*7*11

ganeshie8 (ganeshie8):

\[x \equiv 3 (7\cdot 11)[ \color{Red}{3}] \\+\\5 (5\cdot 11)[ \color{Red}{6}] \\+ \\7 (5\cdot 7)[\color{Red}{ 6}]\\ \pmod{5\cdot 7 \cdot 11}\] simplifies to \[x \equiv 3813 \pmod{5\cdot 7 \cdot 11}\]

ganeshie8 (ganeshie8):

which further reduces to \[x \equiv 348\pmod{5\cdot 7 \cdot 11}\] because 3813 = 9(5*7*11) + 348

OpenStudy (mathmath333):

yes i got that by my weird calculations

ganeshie8 (ganeshie8):

we're good then :) calculations can get more weird than this. Try below classic problems involving chinese remainder thm when free

OpenStudy (mathmath333):

i got this lol \(\large \begin{align} \color{black}{ \dfrac{4}{5}+\dfrac{2}{7}+\dfrac{9}{11}\hspace{.33em}\\~\\ \dfrac{733}{385}\hspace{.33em}\\~\\ \dfrac{348+385}{385}\hspace{.33em}\\~\\ 348\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

oh thnks for the problem exersize

ganeshie8 (ganeshie8):

looks interesting xD i don't get how you worked it though, could u elaborate ?

OpenStudy (mathmath333):

yes wait

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \dfrac{3\times 7\times 11\times 3+5\times 5\times 11\times 6+7\times 5\times 7\times 6}{5\times 7\times11} \hspace{.33em}\\~\\ \implies \dfrac{3\times 7\times 11\times 3}{{5\times 7\times11}}+\dfrac{5\times 5\times 11\times 6}{{5\times 7\times11}}+\dfrac{7\times 5\times 7\times 6}{5\times 7\times11} \hspace{.33em}\\~\\ \implies \dfrac{9}{{5}}+\dfrac{30}{{7}}+\dfrac{42}{11} \hspace{.33em}\\~\\ \implies \dfrac{4}{{5}}+\dfrac{2}{{7}}+\dfrac{9}{11} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

that continues to the previous one

ganeshie8 (ganeshie8):

Very interesting xD but il need to think about it more as i suspect this way of simplification can get us into problems sometimes hmm

OpenStudy (mathmath333):

i dont think it is testified by me \(\large \begin{align} \color{black}{ \implies \dfrac{4}{{5}}+\dfrac{2}{{7}}+\dfrac{9}{11} \hspace{.33em}\\~\\ \implies \dfrac{-1}{{5}}+\dfrac{2}{{7}}+\dfrac{-2}{11} \hspace{.33em}\\~\\ \implies \dfrac{-37}{385} \hspace{.33em}\\~\\ \implies \dfrac{385-37}{385} \hspace{.33em}\\~\\ \implies \dfrac{348}{385} \hspace{.33em}\\~\\ \implies 348 }\end{align}\) see it is also working with negatives

ganeshie8 (ganeshie8):

yeah i dont see anything wrong, looks cool :D

OpenStudy (anonymous):

simply use chines remainder theorem